OpenStudy (anonymous):

-7x + 3 = -15/2 please help me

OpenStudy (whpalmer4):

\[-7x+3 = -\frac{15}{2}\] Why don't you start by multiplying the whole thing by 2? That gives \[-7x*2 + 3*2 = -\frac{15}{\cancel{2}}\cancel{*2}\]\[-14x+6 = -15\]Next, subtract 6 from both sides, and divide both sides by the number in front of \(x\) after you've done that. That should give you the value of \(x\).